OpenStudy (anonymous):

You role a standard number cube. find P number grater then 1 6/5 5/6 1/6 1

OpenStudy (p0sitr0n):

P(x>1) = 1- P(x<=1) = 1- 1/6= 5/6

OpenStudy (anonymous):

thanks